# Towards Human Understanding of Paraphrase Types in ChatGPT

### v1

@@ Abstract
Paraphrases represent a human’s intuitive ability to understand expressions presented in various different ways. Current paraphrase evaluations of language models primarily use binary approaches, offering limited interpretability of specific text changes. Atomic paraphrase types (APT) decompose paraphrases into different linguistic changes and offer a granular view of the flexibility in linguistic expression (e.g., a shift in syntax or vocabulary used). In this study, we assess the human preferences towards ChatGPT in generating English paraphrases with ten APTs and five prompting techniques. We introduce APTY (Atomic Paraphrase TYpes), a dataset of 500 sentence-level and word-level annotations by 15 annotators. The dataset also provides a human preference ranking of paraphrases with different types that can be used to fine-tune models with RLHF and DPO methods. Our results reveal that ChatGPT can generate simple APTs, such as additions and deletions, but struggle with complex structures (e.g., subordination changes). This study contributes to understanding which aspects of paraphrasing language models have already succeeded at understanding and what remains elusive. In addition, our curated datasets can be used to develop language models with specific linguistic capabilities.
@@ 1 Introduction
Atomic Paraphrase Types (APT) Barrón-Cedeño et al. (2013); Vila et al. (2014) can be used as a new lens with which the linguistic relationship between two paraphrases can be explained. Generating and detecting APTs over binary categorizing paraphrases has multiple advantages Wahle et al. (2023). For example, APTs can pinpoint whether a sentence’s grammatical structure or the used vocabulary has changed between potential plagiarism cases Alvi et al. (2021). Understanding how language models understand this variation in linguistic expression gives us insights into how their understanding of language differs from that of humans. It also explains in which language aspects models are proficient, where challenges remain, and how we can make models more robust to a wide array of paraphrase characteristics (e.g., syntactical and lexical changes). There are many ways in which two paraphrases can differ. Consider the following example:
@@
So far, it has been largely unknown how well models generate or detect paraphrases with specific APTs. In this work, we asked 15 humans to annotate 500 APT generations with various properties such as the perceived difficulty of generation, the model’s success at generating a certain type and the reasons behind their failure, its confusion with other types, and the similarity of the APT on a sentence and word level. We publish this dataset as Atomic Paraphrase TYpes Base (APTYBase) [^1] https://github.com/worta/apty. To further contribute to the future of research on APTs, we extended this new dataset by ranking the different APT examples by human preferences that can be used to optimize paraphrase generation models in the future using human preference methods such as RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) called APTYRanked. The whole generation and annotation process is shown in Figure 1 and discussed in Section 3.1.
Our results show that ChatGPT-3.5-turbo-0613 (ChatGPT) is capable of generating Same Polarity Substitution, Semantic-Based Changes, and Change of Order and struggle with generating Inflectional Changes, Synthetic/Analytic Substitutions, and Subordination and Nesting Changes. In general, changes requiring deeper grammatical understanding are difficult to generate. Few-shot and chain-of-thought (CoT) prompting have increased generation success compared to other prompting techniques, especially for Addition/Deletion and Semantic-Based Changes. Surprisingly, humans often rank CoT generations lower than other methods. We also found that the most common error for ChatGPT when applying APT is applying the wrong kind of APT and that morpholexical changes, i.e., changes that arise at the word or morpheme level, are the most commonly wrongly or additionally applied. Lastly, our results suggest that the human-estimated hardness of the generation task does not affect the generation success rate for zero-, one- and few-shot prompts. CoT and a fine-tuned model perform markedly worse for tasks rated as hard. Our main contributions are:
@@
    Analysis of human preferences (Q3) of paraphrase generation and a new dataset APTYRanked with human-ranked paraphrase type generations using best-worst scale which enables training RLHF and DPO methods;
@@
    Correlation of model success rate at generating paraphrases and human perceived difficulty (Q6) of the task.
@@ 2 Related Work
Approaches for paraphrase generation range from rule- and template-based approaches Androutsopoulos and Malakasiotis (2009) to trained transformers generating paraphrased text Wahle et al. (2022). Rule-based methods rely on parsing the original sentence and applying either hand-crafted McKeown (1983) or automatically inferred LIN and PANTEL (2001) rules to transform the text. Recently, paraphrase generation involves deep learning models, especially LLMs. Witteveen and Andrews (2019) use a fine-tuned version of GPT-2 to generate paraphrases and evaluate their semantic similarity. Palivela (2021) fine-tune a T5 model to generate and identify paraphrases. Wahle et al. (2022) explore T5 and GPT-3 regarding qualitative properties of generated paraphrases, access the ability of humans to identify machine-paraphrased textm and suggest that LLMs can generate paraphrases that match human-generated paraphrases in clarity, fluency, and coherence.
As previous paraphrase tasks rely heavily on similarity scores and do not capture the linguistic flexibility of paraphrases, Wahle et al. (2023) proposed two new tasks, i.e., paraphrase type generation and detection, using the ETPC Kovatchev et al. (2018) dataset. Their findings indicate that current LLMs (e.g., ChatGPT) perform well when generating paraphrases with generic semantic similarity but struggle to generate them with fixed APTs. Additionally, models trained with APTs have improved performance in general paraphrase tasks (i.e., without APTs). While little is known yet in full detail on the paraphrastic mechanisms of all paraphrase types, further work revealed that specific types elicit prompt engineering capabilities over various downstream tasks, e.g., polarity for sentiment, or discourse for summarization Wahle et al. (2024).
Although Wahle et al. (2023) has sparked interest in more granular paraphrasing, their work relies only on automatic metrics such as ROUGE Lin (2004), which is limited and lacks the human component in the evaluation process. We explore human preferences in paraphrases by examining whether a paraphrase type was correctly applied by a model, the perceived difficulty of the paraphrasing task, and the kind of errors made by ChatGPT at the time of generation.
@@ 3.1 Paraphrase Type Generation
We sample ten paraphrase pairs for each paraphrase type $l$. We use the first sentence in a paraphrase pair as a base sentence and generate the paraphrase using ChatGPT-3.5-turbo-0613. The prompts are constructed by asking the model to generate a sentence with the same meaning using the APT $l$. We prompt the model using zero-shot, one-shot, few-shot Bragg et al. (2021), chain-of-thought (CoT) Wei et al. (2022), and a fine-tuned model from Wahle et al. (2023). The prompt contains the name of type $l$ and its definition, taken from Barrón-Cedeño et al. (2013) with minimal changes to fit the prompt (see Appendix C for exact prompts used.) Similarly to Bragg et al. (2021), five few-shot examples are given; one instance with added reasoning is provided for CoT prompts. Appendix D details the exact APT definitions.
@@ 4 Experiments
Note that for Same Polarity Substitution, no change in sentence structure or on word level somewhere else in the sentence is necessary, i.e., the change is locally constrained and does not require understanding beyond synonym relations. However, changing smiling to smiled makes it necessary to shift the word’s position, as it is now used as a verb. It needs to be conjugated to match the rest of the sentence, which is in the simple past tense and third person singular and requires the addition of and to keep the sentence correct. The high success rate of 82% for Semantic Changes is surprising as a semantic shift is typically complex. As an example, consider: Original: “Blair himself took the reins of the Labour Party(a).” Paraphrase: “Blair himself became leader of the Labour Party(a).” However, the lack of correspondence also gives the models the most leeway, as the whole grammatical structure can be exchanged. Additionally, the change often involves idioms and other common turns of phrases, which might be seen often in training data.
Structure-based changes present further difficulty for the model. These changes require a deep grammatical understanding of the text. More complex paraphrase scenarios seem more difficult to understand for LLMs, although based on performance in the general paraphrasing task, they should be capable of handling them. In the general paraphrase generation task without APTs, models might seem to generate diverse paraphrases but repeatedly use a small selection of APTs that convince annotators of diversity as a form of reward hacking Casper et al. (2023). One explanation could be that LLMs are more familiar with some paraphrase types than others, i.e., paraphrasing is also biased towards certain types in the training data and fine-tuning steps Zhou et al. (2022a). Humans use different paraphrase types in varying degrees which is reflected in human-created training data; this might explain this observation. Without diversity of paraphrase types in training, models might be restricted in linguistic diversity, even in cases where underrepresented paraphrase types would be advantageous, similar to how it hinders them in detection tasks Zhou et al. (2022a).
@@
Besides whether a change is applied correctly, the resulting paraphrases have varying degrees of quality. Humans might prefer one paraphrase over another, although they have applied the same type (e.g., some grammatical structures are easier to grasp than others). These preferences can indirectly inform what criteria humans use to judge paraphrases, e.g., at which position a change happens or how creative a change is. We investigate whether different prompt techniques (e.g., CoT, zero-shot) produce paraphrases rated as more or less favorable by humans.
@@
As generations from few-shot prompts are perceived best on average, it suggests that practical examples allow ChatGPT to generate natural-sounding paraphrases more than reasoning. Further, while CoT has one of the highest success rates in generating APTs, it seems to generate paraphrases of lesser quality. One reason for the poor reception of CoT-generated paraphrases might be that more consideration of the formal aspects of a paraphrase leads to less natural outcomes, which are not perceived well by the annotators. The generations of the fine-tuned model were also rated lower than prompting the general model. We noted that the fine-tuned model struggled to generate specific APT changes; this seems to extend to natural-sounding paraphrases with single APTs. That highlights the importance of high-quality datasets tailored to the specific tasks for fine-tuned models. Our preference data can be used in future work to improve model generations via techniques like RLHF or DPO.
@@
Ans. In the previous section, we have seen that ChatGPT often applies different APTs if an error occurs. ChatGPT also applies changes besides the desired one in the case of successful applications. In such instances, we can examine the relationship between the different APTs from the perspective of ChatGPT, assessing whether and how they correlate. For these cases, the annotators indicated which kind of APT groups these erroneous or additional changes belong to. We used four major groups for annotators to indicate confusion: morpho-lexical, semantic, structural, and other changes; the definitions can be found in the Section E.4.
@@
For morpho-lexical changes, the additional changes are evenly spread and happen comparatively rarely. If ChatGPT makes an erroneous change, it is often a different morpho-lexical change. This is also the case for the other groups, e.g., the most common type of wrong application concerns the model making morpho-lexical changes instead of the requested changes. These are also probably the most straightforward changes to make without changing the meaning of the sentence and also the most common change, e.g., the most common types in the ETPC dataset are morpho-lexical changes. For Semantic Changes, there is an association with a high proportion of additional/erroneous morpho-lexical changes and structural changes, but not conversely. For example, ChatGPT often mistakes morpho-lexical changes for semantic ones but not vice versa. As Semantic Changes include all changes on a semantic level without clear mapping between the linguistic components, models often introduce other APT simultaneously. Generally, the boundaries between different types of the same group seem larger than between different groups. So, a model focused on one APT group, e.g., structural changes, should be more reliable than one handling multiple APT groups. Otherwise, models often introduce other changes. One reason might be that models rarely see isolated changes in the training data (i.e., sentences that only differ by a single APT), leading to a poor understanding and associating them with other changes. We also found that the confusion correlates with the difficulty of the paraphrasing task; the more difficult the task is, the more likely it is that other changes are made either in addition or erroneously. We provide more detailed investigations in the additional research questions in Section B.1.
Q6. How does ChatGPT perform on examples perceived as easy or hard by humans?
Ans. Applying different APTs can be challenging for humans, depending on the precise APT and the base sentence. We explore if the human perceived difficulty of paraphrases also relates to the performance of ChatGPT; that is, are paraphrasing tasks difficult for humans also difficult for ChatGPT?

### v2

# Towards Human Understanding of Paraphrase Types in Large Language Models

@@ Abstract
Paraphrases represent a human’s intuitive ability to understand expressions presented in various different ways. Current paraphrase evaluations of language models primarily use binary approaches, offering limited interpretability of specific text changes. Atomic paraphrase types (APT) decompose paraphrases into different linguistic changes and offer a granular view of the flexibility in linguistic expression (e.g., a shift in syntax or vocabulary used). In this study, we assess the human preferences towards ChatGPT in generating English paraphrases with ten APTs and five prompting techniques. We introduce APTY (Atomic Paraphrase TYpes), a dataset of 800 sentence-level and word-level annotations by 15 annotators. The dataset also provides a human preference ranking of paraphrases with different types that can be used to fine-tune models with RLHF and DPO methods. Our results reveal that ChatGPT and a DPO-trained LLama 7B model can generate simple APTs, such as additions and deletions, but struggle with complex structures (e.g., subordination changes). This study contributes to understanding which aspects of paraphrasing language models have already succeeded at understanding and what remains elusive. In addition, we show how our curated datasets can be used to develop language models with specific linguistic capabilities.
Towards Human Understanding of Paraphrase Types in Large Language Models
Dominik Meier1,2,3, Jan Philip Wahle1, Terry Ruas1, Bela Gipp1 1University of Göttingen, Germany 2LKA NRW, Germany 3meier@gipplab.org
@@ 1 Introduction
Atomic Paraphrase Types (APT) Barrón-Cedeño et al. (2013); Vila et al. (2014) can be used as a new lens through which the linguistic relationship between two paraphrases can be explained. Generating and detecting APTs over binary categorizing paraphrases has multiple advantages Wahle et al. (2023). For example, APTs can pinpoint whether a sentence’s grammatical structure or the used vocabulary has changed between potential plagiarism cases Alvi et al. (2021). Understanding how language models understand this variation in linguistic expression gives us insights into how their understanding of language differs from that of humans. It also explains in which language aspects models are proficient, where challenges remain, and how we can make models more robust to a wide array of paraphrase characteristics (e.g., syntactical and lexical changes). There are many ways in which two paraphrases can differ. Consider the following example:
@@
So far, how well models generate or detect paraphrases with specific APTs has been largely unknown. In this work, we asked 15 humans to annotate 800 APT generations with various properties such as the perceived difficulty of generation, the model’s success at generating a certain type and the reasons behind their failure, its confusion with other types, and the similarity of the APT on a sentence and word level. We publish this dataset as Atomic Paraphrase TYpes Base (APTYBase) [^1] https://github.com/worta/apty. To further contribute to the future of research on APTs, we extended this new dataset by ranking the different APT examples by human preferences that can be used to optimize paraphrase generation models in the future using human preference methods such as RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) called APTYRanked. We demonstrate the usefulness of the preference data by using DPO to train a Llama 2 7B Touvron et al. (2023) based model and compare the generation performance to a supervised fine-tuned and base version. The whole generation and annotation process is shown in Figure 1 and discussed in Section 3.1.
Our results show that ChatGPT-3.5-turbo-0613 (ChatGPT) is capable of generating Same Polarity Substitution, Semantic-Based Changes, and Change of Order and struggle with generating Inflectional Changes, Synthetic/Analytic Substitutions, and Subordination and Nesting Changes. In general, changes requiring deeper grammatical understanding are difficult to generate. Few-shot and chain-of-thought (CoT) prompting have increased generation success compared to other prompting techniques, especially for Addition/Deletion and Semantic-Based Changes. Surprisingly, humans often rank CoT generations lower than other methods. We also found that the most common error for ChatGPT when applying APT is applying the wrong kind of APT and that morpholexical changes, i.e., changes that arise at the word or morpheme level, are the most commonly wrongly or additionally applied. Lastly, the DPO-trained LLama outperforms the supervised fine-tuned and base Llama models by a wide margin. Our main contributions are:
@@
    Analysis of human preferences (Q3) of paraphrase generation and a new dataset APTYRanked with human-ranked paraphrase type generations using best-worst scale
@@
    Evaluation of DPO training (Q6) with the obtained human preferences based on the LLama 2 7B model family with 300 annotations
@@ 2 Related Work
Approaches for paraphrase generation range from rule- and template-based approaches Androutsopoulos and Malakasiotis (2009) to trained transformers generating paraphrased text Wahle et al. (2022). Rule-based methods rely on parsing the original sentence and applying either hand-crafted McKeown (1983) or automatically inferred Lin and Pantel (2001) rules to transform the text. Recently, paraphrase generation involves deep learning models, especially LLMs Zhou and Bhat (2021). Witteveen and Andrews (2019) use a fine-tuned version of GPT-2 to generate paraphrases and evaluate their semantic similarity. Palivela (2021) fine-tune a T5 model to generate and identify paraphrases. Wahle et al. (2022) explore T5 and GPT-3 regarding qualitative properties of generated paraphrases, access the ability of humans to identify machine-paraphrased text, and suggest that LLMs can generate paraphrases that match human-generated paraphrases in clarity, fluency, and coherence.
As previous paraphrase tasks rely heavily on similarity scores and do not capture the linguistic flexibility of paraphrases, Wahle et al. (2023) proposed two new tasks, i.e., paraphrase type generation and detection, using the ETPC Kovatchev et al. (2018) dataset. Their findings indicate that current LLMs (e.g., ChatGPT) perform well when generating paraphrases with generic semantic similarity but struggle to generate them with fixed APTs. Additionally, models trained with APTs have improved performance in general paraphrase tasks (i.e., without APTs). While little is known yet in full detail on the paraphrastic mechanisms of all paraphrase types, further work revealed that specific types elicit prompt engineering capabilities over various downstream tasks, e.g., polarity for sentiment, or discourse for summarization Wahle et al. (2024). An alternative approach to APT to have more control over linguistic properties is generating paraphrases based on syntactic templates Huang and Chang (2021). However, changes on a template level are too specific in a generation or detection task for humans to specify desired concepts.
Although Wahle et al. (2023) has sparked interest in more granular paraphrasing, their work relies only on automatic metrics such as BLEU Papineni et al. (2002) or ROUGE Lin (2004), which is limited and lacks the human component in the evaluation process. We explore human preferences in paraphrases by examining whether a paraphrase type was correctly applied by a model, the kind of errors made by ChatGPT at the time of generation, and by ranking paraphrases according to human judgement.
@@ 3.1 Paraphrase Type Generation
We sample ten paraphrase pairs for each paraphrase type $l$. We use the first sentence in a paraphrase pair as a base sentence and generate the paraphrase using ChatGPT-3.5-turbo-0613. The prompts are constructed by asking the model to generate a sentence with the same meaning using the APT $l$. We prompt the model using zero-shot, one-shot, few-shot Bragg et al. (2021), chain-of-thought (CoT) Wei et al. (2022), and a fine-tuned model from Wahle et al. (2023), which is based on ChatGPT-3.5-turbo-0613 as well. The prompt contains the name of type $l$ and its definition, taken from Barrón-Cedeño et al. (2013) with minimal changes to fit the prompt (see Appendix C for exact prompts used.) Similarly to Bragg et al. (2021), five few-shot examples are given; one instance with added reasoning is provided for CoT prompts. Appendix D details the exact APT definitions. We use the same base sentences for evaluating the DPO-tuned LLama model and follow the prompt template for the fine-tuned model. We use Llama 2 7B, a version of LLama 2 7B fine-tuned on the ETPC dataset from Wahle et al. (2023), and a DPO-trained Llama model trained by us on the ranked data from the second phase of the annotation and based on the fine-tuned model.
@@ 3.2 Annotation
For the evaluation of the DPO-tuned Llama model, two annotators performed the first phase again, but each annotated every one of the 300 (3 models x 10 APT x 10) generations.
@@ 4 Experiments
Summary Change of Order, Semantic Changes, and Same Polarity Substitution can be generated with decent success rates, while more complicated changes, particularly those involving deeper grammatical changes at multiple points like Derivational Changes pose a problem.
@@
Note that for Same Polarity Substitution, no change in sentence structure or on word level somewhere else in the sentence is necessary, i.e., the change is locally constrained and does not require understanding beyond synonym relations. However, changing smiling to smiled makes it necessary to shift the word’s position, as it is now used as a verb. It needs to be conjugated to match the rest of the sentence, which is in the simple past tense and third person singular and requires the addition of and to keep the sentence correct. The high success rate of 82% for Semantic Changes is surprising as a semantic shift is typically complex. But the change also often involves idioms and other common turns of phrases, which might be seen often in training data. Structure-based changes present further difficulty for the model. These changes require a deep grammatical understanding of the text.
In the general paraphrase generation task without APTs, models might seem to generate diverse paraphrases but repeatedly use a small selection of APTs that convince annotators of diversity as a form of reward hacking Casper et al. (2023). One explanation could be that LLMs are more familiar with some paraphrase types than others, i.e., paraphrasing is also biased towards certain types in the training data and fine-tuning steps Zhou et al. (2022a). Without diversity of paraphrase types in training, models might be restricted in linguistic diversity, even in cases where underrepresented paraphrase types would be advantageous, similar to how it hinders them in detection tasks Zhou et al. (2022a).
@@
Summary CoT is the most successful prompting strategy, while one-shot prompting led to the highest error rate.
@@
Summary Annotators favor the generations from few-shot prompting the most, while generations from CoT and the Fine-Tuned model are generally preferred less.
Ans. Besides whether a change is applied correctly, the resulting paraphrases have varying degrees of quality. Humans might prefer one paraphrase over another, although they have applied the same type (e.g., some grammatical structures are easier to grasp than others). These preferences can indirectly inform what criteria humans use to judge paraphrases, e.g., at which position a change happens or how creative a change is. We investigate whether different prompt techniques (e.g., CoT, zero-shot) produce paraphrases rated as more or less favorable by humans.
@@
As generations from few-shot prompts are perceived best on average, it suggests that practical examples allow ChatGPT to generate natural-sounding paraphrases more than reasoning. Further, while CoT has one of the highest success rates in generating APTs, it seems to generate paraphrases of lesser quality. One reason for the poor reception of CoT-generated paraphrases might be that more consideration of the formal aspects of a paraphrase leads to less natural outcomes, which are not perceived well by the annotators. The generations of the fine-tuned model were also rated lower than prompting the general model. We noted that the fine-tuned model struggled to generate specific APT changes; this seems to extend to natural-sounding paraphrases with single APTs. That highlights the importance of high-quality datasets tailored to the specific tasks for fine-tuned models. Our preference data can be used to improve model generations via techniques like RLHF or DPO.
@@
Summary The most common type of wrong application concerns the model making morpho-lexical changes instead of the requested changes and the boundaries between different types of the same group seem larger than between different groups. Applying the wrong APT is the most frequent reason for error by far.
@@
Summary We see intra-group errors less often than inter-group errors and morpho-lexical changes are applied erroneously most often.
Ans. The previous answer shows that ChatGPT often applies different APTs if an error occurs. ChatGPT also applies changes besides the desired one in the case of successful applications. In such instances, we can examine the relationship between the different APTs from the perspective of ChatGPT, assessing whether and how they correlate. For these cases, the annotators indicated which kind of APT groups these erroneous or additional changes belong to. We used four major groups for annotators to indicate confusion: morpho-lexical, semantic, structural, and other changes; the definitions can be found in the Section E.4.
@@
For morpho-lexical changes, the additional changes are evenly spread and happen comparatively rarely. If ChatGPT makes an erroneous change, it is often a different morpho-lexical change. This is also the case for the other groups, e.g., the most common type of wrong application concerns the model making morpho-lexical changes instead of the requested changes. These are probably the most straightforward changes to make without changing the meaning of the sentence and also the most common change, e.g., the most common types in the ETPC dataset are morpho-lexical changes. Generally, the boundaries between different types of the same group seem larger than between different groups. One reason might be that models rarely see isolated changes in the training data (i.e., sentences that only differ by a single APT), leading to a poor understanding and associating them with other changes. We also found that the confusion correlates with the difficulty of the paraphrasing task; the more difficult the task is, the more likely it is that other changes are made either in addition or erroneously. We provide more detailed investigations in the additional research questions in Section B.1.
Q6. Can the human preference data improve the generation of APTs?
Summary Training a model with the collected preference data improves the performance of the model. We can see similar trends in LLama as we see in ChatGPT.
Ans. We used the collected preferences to train a LLama3 7B model Dubey et al. (2024), which was fine-tuned on the ETPC dataset from Wahle et al. (2023), via DPO Rafailov et al. (2023). We compare the results in generation success across LLama models (base, fine-tuned, DPO-trained); the results for each APT are shown in Figure 5. Across all generations, DPO performs best with a rate of 34% of successful APT applications, followed by the base model with 14% and the fine-tuned model with 11%. We can also see that, while success rates are generally much lower than with the much larger model, the trends from ChatGPT also hold across the tested LLama 7B variants, i.e., success rates across all variants are best for types where changes are local and where the model has freedom, like Semantic Changes (37%), and Addition/Deletion (43%). At the same time, they are worse for changes that require a deeper grammatical understanding like Derivational Changes (8%). We see that the DPO training improves generation success by a large margin, and the collected preference data improves the understanding of the model significantly. The success rates for difficult changes like Synthentic/Analytic substitution (10% to 40%) or Subordination and Nesting Changes (10% to 50%), e.g., changes which we have seen ChatGPT struggles to perform as well, received a big boost via DPO training, suggesting an increased understanding of the model of these concepts. Surprisingly, fine-tuning on the ETPC dataset does not improve generation success in this task. One explanation might be that in ETPC multiple changes are performed at once, and the model does not learn the specific mechanics of a single change due to the notice introduced by other changes.
Figure 5: The success rate in generating a specific APT for different LLama 7B Models